HIGH RESOLUTION LASER SPECTROSCOPY OF THE A0^{+} − X0^{+} SYSTEM OF YbS

\underline{TODD C. MELVILLE}, JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, NS, Canada B3H 4J3; COLAN LINTON, Physics Department, University of New Brunswick, P.O. Box 4400, Fredericton, NB, Canada E3B 5A3.

The first spectroscopic observation of YbS has been made using visible laser absorption spectroscopy. The A0^{+} − X0^{+} electronic transition of ^{172}Yb^{32}S, ^{174}Yb^{32}S and ^{176}Yb^{32}S has been recorded using the output of a Coherent 699-29 ring dye laser with selective detection of fluorescence. Spectra of the 0-0, 0-1, 1-1, 1-2, 2-2 and 2-3 bands were recorded with a measurement precision of approximately 0.003 cm^{−1}, and from perturbations observed in the structure of the 2-2 and 2-3 bands, it was concluded that a level crossing occurs in the A0^{+} v=2 level. In total, over 1500 line positions have been measured, assigned and employed in least- squares fits of the molecular parameters. The deperturbation and rotational analysis of the A − X system will be discussed.